OpenStudy (anonymous):

please help!! 1.5^lnx=10^0.5 find the solution and round to the nearest thousandth

OpenStudy (mertsj):

ln(x)log(1.5)=.5log10

OpenStudy (mertsj):

Divide both sides by log1.5

OpenStudy (mertsj):

e raised to that power will be x

OpenStudy (anonymous):

1.5^lnx=10^0.5 take natural log on both sides gives lnx ln1.5=0.5ln10 ln x=(0.5ln 10)/ln1.5=2.839 e^lnx=e^2.839 x=17.106

OpenStudy (anonymous):

Thanks!!!!!

OpenStudy (mertsj):

\[\ln x \log_{10}1.5=.5\log_{10}10 \]

OpenStudy (mertsj):

\[\ln x=\frac{.5}{\log_{10}1.5} \]

OpenStudy (mertsj):

\[\ln x=2.8394\]

OpenStudy (mertsj):

\[e ^{2.8394}=x=17.106\]

OpenStudy (anonymous):

thank you also